The solution to this equation 6y+7=5-y
Answer by farohw(175)


For 6y + 7 = 5 - y you want to group like terms then calculate as follows:

6y + y + 7 -5 = 0

7y + 2 = 0

7y = -2

Divide both sides by 7 to get y = -2/7

You can also write it as

6y + 7 = 5 - y ---> 6y + y = 5 - 7

7y = -2

y = -2/7

Either way you will get the same answer :)

Cheers!